OpenStudy (anonymous):

Find the monthly payment for the loan. (Round your answer to the nearest cent.) A $120,000 home bought with a 20% down payment and the balance financed for 30 years at 10.5%

OpenStudy (wolf1728):

The formula for this is pretty intimidating. I answered a question like this last night so I'll try again. The loan is for 120,000 then after the 20% down payment of 24,000, the loan is for 96,000.00 for 30 years at 10.5% We need the rate divided by 1200 so that equals 0.00875 (monthly interest) number of months = 360 monthly payment = {.00875 + .00875 / [((1+.00875)^360)-1]} * 96,000 monthly payment = (.00875 + .00875 /22.0185086616) * 96,000 monthly payment = (.00875 + 0.000397392944930708) * 96,000 monthly payment = (0.00914739294493071) * 96,000 monthly payment = 878.1497227133 monthly payment = 878.15 rounded (I checked that with a loan calculator - it is correct).

OpenStudy (anonymous):

Ok thank you

OpenStudy (wolf1728):

u r welcome. (as I said that formula is very intimidating.)

OpenStudy (anonymous):

yes it is I have another one I'm working on

OpenStudy (wolf1728):

Okay - let me know - I'll stick around.